Suppose there are only two individuals in Winterwonderland, Snow White and Thor. Snow White’s demand curve for snow removal is given by qsw =50−2p For Thor, the demand curve for snow removal is given by qthor = 100 − p

a. Suppose snow removal is a pure public good; that is, once it is done, everyone benefits from it. What would be the optimal level of this activity if it could be produced at a constant marginal cost of $80 per mile?

b. If snow removal were left to the private market, how much might be produced?
Answer #1

a. qsw =50−2p

2p= 50 - qsw

P= 25 - 0.5qsw

qthor = 100 − p

P= 100 - qthor

Aggregate demand = P= 25 - 0.5q + 100 - q

P= 125 - 1.5q

P= MC at optimum

125 - 1.5q = 80

q= 30

b) TR= PQ = 125q- 1.5q^2

MR= 125 - 3q

MR = MC

125 - 3q = 80

q= 15

P= 125 - 1.5(15)= 102.5